# Derivative-enhanced Deep Operator Network

### v1

@@ 3.3 Dimension Reduction
and its right hand side through the matrix-vector multiplication in Equation 12 (See  Section A.2 for details). Similar to the KLE case, let $\Psi^{\text{in}}$ denote the nodal values of $r$ dominant eigenfunctions. Since the eigenvectors $\bm{\psi}_{i}^{\text{in}}$ are $C^{-1}$-orthogonal, the reduced representation of $\bm{m}$ can be computed as $\widetilde{\bm{m}}=(\Psi^{\text{in}})^{T}C^{-1}\bm{m}$.
For the sake of scalability, we use the double pass randomized algorithm [18] in hIPPYlib (an open-source inverse problem python library) to solve the generalized eigenproblems Equation 7 and Equation 10.
3.4 Computation of the action of $\mathcal{H}$
By Equation 9, we use $N_{\text{grad}}$ samples $\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\text{grad}}}$ to compute the Monte Carlo estimate of the action of operator $\mathcal{H}$ on any function $\psi\in V_{h}^{\text{in}}$, that is,
Each summation term in Equation 13 can be obtained by computing $p:=du(m;\psi)$ followed by $d^{*}u(m;p)$.
@@ 3.4 Computation of the action of $\mathcal{H}$
We first show how to compute $p$ given the PDE residual $\mathcal{R}(m,u)=0$. Since $u$ is uniquely determined by $m$, we can write $u=u(m)$ so that $\mathcal{R}(m,u(m))\equiv 0$ holds for any $m\in V_{h}^{\text{in}}$. Thus, for any $\varepsilon>0$ and $\psi\in V_{h}^{\text{in}}$, we have
@@
where $v$ is a test function in $V_{h}^{\text{out}}$.
@@
Let $q:=(\partial_{u}\mathcal{R})^{-*}p$. We can solve for $p$ via the weak form
$\langle(\partial_{u}\mathcal{R})^{*}q,v\rangle=\langle p,v\rangle,$ (18)
where $v$ is a test function in $V_{h}^{\text{out}}$.
By Equation 17, we have $w=-(\partial_{m}\mathcal{R})^{*}q$. For any test function $v\in V_{h}^{\text{in}}$, it holds that
@@
When using the double pass randomized algorithm to obtain the first $r$ eigenpairs in  Equation 10, we need to compute the action of $\mathcal{H}$ on $2(r+s)$ random functions $V_{h}^{\text{in}}$ (e.g., their nodal values are sampled from the standard Gaussian distribution), where $s\in\mathbb{N}^{+}$ is typically a small oversampling parameter, often chosen between $5$ and $20$. To speed up the computation, we first compute the LU factorization of the matrices resulting from the discretization of the linear PDEs in Equation 16 and Equation 18. Then the action of $d^{*}u(m^{(i)};du(m^{(i)};\cdot))$ on these random functions can be efficiently computed via the forward and backward substitution. Furthermore, the computational time can be significantly reduced by parallelizing the computation of the average value in Equation 13 across multiple processors.
4 Experiments
In this section, we present experiment results to demonstrate the derivative-enhanced accuracy and cost of our method on three test PDEs, including both linear and nonlinear, both scalar-valued and vector-valued solutions. Details about the data generation and training can be found in Appendix B.
4.1 Input Probability Measure
In all test cases, we assume that the input functions $m^{(i)}$ are i.i.d. samples drawn from a Gaussian random field with mean function $\bar{m}$ and covariance operator $(\delta I-\gamma\Delta)^{-\alpha}$, with $\alpha=2$ so the covariance operator is of trace class. It is worth noting that the parameters $\gamma$ and $\delta$ jointly control the marginal variance and correlation length, for which we take values to have large variation of the samples so that it is difficult for the neural network approximation with relatively large errors. We use a scalable sampling algorithm in hIPPYlib [18] to generate the samples.
4.2 Governing Equations
We consider a linear and scalar-valued Darcy flow equation, a nonlinear vector-valued hyperelasticity equation with one state (displacement), and a nonlinear vector-valued Navier–Stokes equations with multiple states (velocity and pressure). Specifically, we learn the map from (the logarithm of) the diffusion coefficient field to the solution of the Darcy flow equation with zero Dirichlet boundary condition. For the hyperelasticity equation, we consider an experimental scenario where a square of hyperelasticity material is secured along its left edge while a fixed upward-right force is applied to its right edge [2]. Our goal is to learn the map between (the logarithm of) the material’s Young’s modulus and its displacement. Lastly, we consider the Navier–Stokes equations that describes viscous, incompressible fluid flow. In particular, we consider the lid-driven cavity case where a square cavity consisting of three rigid walls with no-slip conditions and a lid moving with a tangential unit velocity. We consider that the uncertainty comes from the viscosity term and aim to predict the velocity field.
Darcy flow equation.
The strong form of the PDE can be written as
$\begin{cases}-\nabla\cdot(e^{m}\nabla u)=1,\quad x\in\Omega,\\ u(x)=0,\quad x\in\partial\Omega.\end{cases}$
Here, we assume that the randomness arises from the viscosity term $e^{m}$. For demonstration, we choose $\bar{m}=0$, $\delta=1.0$ and $\gamma=0.1$. Note that the PDE is linear but the solution operator is not. See Figure 1 for the visualization for input-output pairs of the Darcy flow equation.
Figure 1: Visualization of input-output pairs of Darcy flow equation.
Hyperelasticity equation.
$d\widetilde{W}(u;v)=0,$
holds for any test function $v$ in the state space. See Figure 2 for the visualization for input-output pairs of the hyperelasticity equation.
Figure 2: Visualization of input-output pairs of hyperelasticity equation.
Navier–Stokes equation.
where $\Gamma_{\text{top}}$ and $\Gamma$ denote the left and whole boundary of the cavity domain $\Omega$, respectively. Here, we assume that the randomness arises from the viscosity term $e^{m}$. Our goal is learn the operator that maps the parameter $m$ to the velocity $u$. For demonstration, we choose $\bar{m}=6.908$ ($e^{\bar{m}}\approx 10^{3}$), $\delta=0.6$, and $\gamma=0.06$. See Figure 3 for the visualization for input-output pairs of the Navier–Stokes equations.
Figure 3: Visualization of input-output pairs of Navier–Stokes equations.
4.3 Evaluation Metrics
We use the following three metrics to evaluate the performance of different methods. For $u(m)(x)$ approximations, we compute the relative error in the $L^{2}(\Omega)$ norm and $H^{1}(\Omega)$ norm on the test dataset $\mathcal{D}_{\text{test}}=\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\text{test}}}$, that is,
where
For $du(m;\cdot)(x)$ approximations, we compute the relative error in the Frobenius norm on $\mathcal{D}_{\text{test}}$ along random directions $\omega=\{\omega_{i}\}_{i=1}^{N_{\text{dir}}}$, that is,
$\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{m^{(i)}\in\mathcal{D}_{\text{test}}}\frac{\|d\hat{u}(m^{(i)};\omega)(x)-du(m^{(i)};\omega)(x)\|_{F}}{\|du(m^{(i)};\omega)(x)\|_{F}},$
where $\omega_{i}$ are samples drawn from the same distribution as $m$.
4.4 Main Results
We observe consistent performance in terms of enhanced approximation accuracy for the three different equations. Here we only present a part of the results for the Navier–Stokes equations and leave all other results for other equations, including also hyperparameter tuning in the training of the neural networks, the architecture of different neural networks, the configuration of data generations, the full set of test errors, visualization of the KLE and ASM basis functions, and visualization of the prediction of both the solution and derivatives in Appendix B. We show the tradeoff in terms of prediction error and computational cost of different methods. We find that the advantages of our method are especially evident for the more complex and challenging Navier–Stokes equations.
Test errors.
In Figure 4, we show a comparison of different methods for predicting the velocity field in the $y$ direction and its derivative with respect to parameter in the direction of $\{\omega_{i}\}_{i=1}^{N_{\text{dir}}}$ with $N_{\text{dir}}=128$. We observe that the reduced basis DeepONet, even without using derivative labels for training, has a prediction error comparable to that of the vanilla DeepONet. This suggests that the chosen number of reduced bases $(r=16)$ is sufficiently large to predict a high-quality output function.
We also find that incorporating two additional derivative losses (denoted as ‘dm’ loss and ‘dx’ loss) into the training of the reduced basis DeepONet significantly lower the errors (to between one-thirds and one-half of their original values in predicting the velocity field in the $y$ direction) in predicting both the solution $u$ and its directional derivative with respect to the input function $m$. See Figure 5 and Figure 6 for a visualization comparing the prediction of velocity-y and its directional derivative, where the DE-DeepONet with ASM+dm+dx using the two derivative labels provides a more accurate prediction.
In addition, we conduct experiments for comparing the results of adding solely the ‘dm’ loss versus adding both ‘dm’ and ‘dx’ losses. The details are provided in Section B.3. We observe that incorporating ‘dm’ loss makes a dominant contribution to error reduction, while incorporating ‘dx’ loss has a relatively smaller effect on reducing the errors.
Refer to caption Refer to caption Refer to caption
Figure 4: Mean relative errors over 5 trials with a varying number of training samples for the Navier–Stokes equations using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of the velocity-y. Right: Relative error in the Frobenius (Fro) norm for the prediction of the derivative of the velocity-y with respect to the parameter in the direction of the 128 samples drawn from the same distribution as parameter.
Figure 5: Comparison of the predictions of velocity-y with (16, 64, 256, 1024) training samples using different methods. Top: Vanilla DeepONet; Middle: ASM; Bottom: DE-DeepONet using ASM + dm + dx.
Figure 6: Comparison of the predictions of directional derivative $du_{2}(m;\omega_{1})$ with (16, 64, 256, 1024) training samples using different methods. Top: Vanilla DeepONet; Middle: ASM; Bottom: DE-DeepONet using ASM + dm + dx.
Computational cost.
We use MPI and the finite element library FEniCS [10] to distribute the computational load of offline data generation to 64 CPUs for all the PDEs considered in this work. See Table 1 for the wall clock time for generating the samples as Gaussian random fields (GRFs), solving the PDEs, computing the $r=16$ reduced basis functions (KLE or ASM) corresponding to the $16$ largest eigenvalues, generating the derivative labels dm and dx with respect to the parameter and space, respectively. Note that generating the derivative labels is more expensive than generating the PDE solutions, but not by large, especially for the more complex and expensive Navier–Stokes equations.
Table 1: Wall clock time (in seconds) for the data generation on 2 X AMD EPYC 7543 32-Core Processors
Dataset Process
GRFs PDEs KLE ASM dm labels dx labels
$(N_{\text{all}}=2000)$ $(N_{\text{all}}=2000)$ $(r=16)$ $(r=16)$ $(N_{\text{all}}=2000,r=16)$ $(N_{\text{all}}=2000)$
($64$ CPUs) ($64$ CPUs) ($1$ CPU) ($16$ CPUs) ($64$ CPUs) ($64$ CPUs)
Darcy flow $1.1$ $1.4$ $0.4$ $1.1$ $8.3$ $4.6$
Hyperelasticity $1.1$ $9.7$ $0.4$ $1.4$ $19.5$ $8.8$
Navier–Stokes $1.9$ $99.1$ $1.3$ $9.7$ $125.5$ $45.8$
We present comparisons of the computational cost of model training using different methods with two different number of training samples ($N_{\text{train}}=64$, corresponding to a limited training dataset setting, and $N_{\text{train}}=1024$, corresponding to a sufficient training dateset setting) in Table 2. We find that incorporating derivative losses leads to larger computational cost as expected. However, when the training data are limited, the increased computation cost is compensated for a significant reduction of errors. We note that there are potential ways to reduce the cost, e.g., by training the model with additional derivative losses only during the later stage of training, or by using fewer points for computing the derivative losses in each iteration. Additionally, the number of trainable parameters in the neural network is shown in Table 3. We observe that the reduced basis DeepONet has an order of magnitude fewer parameters than the vanilla DeepONet. Thanks to the dimension reduction in the input parameter, we are able to efficiently compute and store the derivatives as well as train the reduced basis DeepONet.
Table 2: Wall clock time (in seconds per epoch) for the neural network training on a single NVIDIA RTX A6000 GPU
Dataset Method ($N_{\text{train}}=64$, $N_{\text{train}}=1024$)
Vanilla Reduced Reduced + dm Reduced + dm + dx
Darcy flow $(0.04,0.64)$ $(0.03,0.50)$ $(0.09,1.58)$ $(0.14,2.39)$
Hyperelasticity $(0.04,0.73)$ $(0.03,0.51)$ $(0.14,2.32)$ $(0.18,3.02)$
Navier–Stokes $(0.04,0.75)$ $(0.04,0.64)$ $(0.15,2.59)$ $(0.22,3.60)$
Table 3: Number of trainable parameters in the neural network
Dataset Neural network size
Vanilla Reduced
Darcy flow $3,420,417$ $101,633$
Hyperelasticity $3,716,482$ $249,858$
Navier–Stokes $3,716,482$ $496,770$
5 Discussion
In this work, we proposed a new neural operator–Derivative-enhanced Deep Operator Network (DE-DeepONet). Compared to vanilla DeepONet, DE-DeepONet receives a reduced representation of input function and has two additional loss terms, which are for the supervised learning of the derivative of the output with respect to the inputs of both the branch net and the trunk net. Our experiments demonstrate that adding these two loss terms to the loss function greatly enhances the accuracy of both function approximations and derivative approximations, especially when the training data are limited.
We observe that when the training data are large enough, the benefits brought by adding derivative loss terms become less significant (that is, with more and more training data, we expect a relatively smaller increase in marginal accuracy and a relatively larger increase in marginal computational cost). However, in computational science and engineering, unlike other machine learning fields such as computer vision, training data in complex and nonlinear physical system applications are often scarce due to the high costs of computational simulations. We demonstrated that the advantage of our method becomes more apparent for the more complex Navier–Stokes equations which require more iterations in the nonlinear solves than the hyperelasticity equation.
Appendix A Proofs
A.1 Proof of the Equivalence of Equation 6 and Equation 7

### v2

@@ 3.3 Dimension reduction
and its right hand side through the matrix-vector multiplication in Equation 12 (See  Section A.3 for details). Similar to the KLE case, let $\Psi^{\text{in}}$ denote the nodal values of $r$ dominant eigenfunctions. Since the eigenvectors $\bm{\psi}_{i}^{\text{in}}$ are $C^{-1}$-orthogonal, the reduced representation of $\bm{m}$ can be computed as $\widetilde{\bm{m}}=(\Psi^{\text{in}})^{T}C^{-1}\bm{m}$. We use a scalable double pass randomized algorithm [28] implemented in hIPPYlib to solve the generalized eigenproblems Equation 7 and Equation 10.
To this end, we present the computation of the derivative label and the action of $\mathcal{H}$ as follows.
Theorem 1.
Suppose the PDE in the general form of (1) is well-posed with a unique solution map from the input function $m\in V^{\text{in}}$ to the output function $u\in V^{\text{out}}$ with dual $(V^{\text{out}})^{\prime}$. Suppose the PDE operator $\mathcal{R}:V^{\text{in}}\times V^{\text{out}}\to(V^{\text{out}})^{\prime}$ is differentiable with derivatives $\partial_{m}\mathcal{R}:V^{\text{in}}\to(V^{\text{out}})^{\prime}$ and $\partial_{u}\mathcal{R}:V^{\text{out}}\to(V^{\text{out}})^{\prime}$, and in addition $\partial_{u}\mathcal{R}$ is invertible with invertible adjoint $(\partial_{u}\mathcal{R})^{*}$. Then the directional derivative $p=du(m;\psi)$ for any function $\psi\in V^{\text{in}}$, and an auxillary variable $q$ such that $d^{*}u(m;p)=-(\partial_{m}\mathcal{R})^{*}q$ can be obtained as the solution of the linearized PDEs
$(\partial_{u}\mathcal{R})p+(\partial_{m}\mathcal{R})\psi$ $=0,$ (13)
$(\partial_{u}\mathcal{R})^{*}q$ $=p.$ (14)
Proof sketch. We perturb $m$ with $\varepsilon\psi$ for any small $\varepsilon>0$ and obtain $\mathcal{R}(m+\varepsilon\psi,u(m+\varepsilon\psi))=0$. Using Taylor expansion to expand it to the first order and letting $\varepsilon$ approach 00, we obtain Equation 13, where $p=du(m;\psi)$. Next we compute $d^{*}u(m;p)$. By Equation 13, we have $du(m;\psi)=-(\partial_{u}\mathcal{R})^{-1}(\partial_{m}\mathcal{R})\psi$. Thus, $d^{*}u(m;\psi)=-(\partial_{m}\mathcal{R})^{*}(\partial_{u}\mathcal{R})^{-*}p$. Then we can first solve for $q=(\partial_{u}\mathcal{R})^{-*}p$ and then compute $-(\partial_{m}\mathcal{R})^{*}q$. See Section A.1 for a full proof.
4 Experiments
In this section, we present experiment results to demonstrate the derivative-enhanced accuracy and cost of our method on two test problems of nonlinear vector-valued PDEs in comparison with DeepONet, FNO, and DINO. Details about the data generation and training can be found in Appendix B.
4.1 Input probability measure
In all test cases, we assume that the input functions $m^{(i)}$ are i.i.d. samples drawn from a Gaussian random field with mean function $\bar{m}$ and covariance operator $(\delta I-\gamma\Delta)^{-\alpha}$. We take $\alpha=2$ in two space dimensions so the covariance operator is of trace class [29]. It is worth noting that the parameters $\gamma$ and $\delta$ jointly control the marginal variance and correlation length, for which we take values to have large variation of the input samples that lead to large variation of the output PDE solutions. In such cases, especially when the mapping from the input to output is highly nonlinear as in our test examples, a vanilla neural network approximations tend to result in relatively large errors, particularly with a limited number of training data. To generate samples from this Gaussian random field, we use a scalable (in mesh resolution) sampling algorithm implemented in hIPPYlib [28].
4.2 Governing equations
We consider two nonlinear PDE examples, including a nonlinear vector-valued hyperelasticity equation with one state (displacement), and a nonlinear vector-valued Navier–Stokes equations with multiple states (velocity and pressure). For the hyperelasticity equation, we consider an experimental scenario where a square of hyperelasticity material is secured along its left edge while a fixed upward-right force is applied to its right edge [30]. Our goal is to learn the map between (the logarithm of) the material’s Young’s modulus and its displacement. We also consider the Navier–Stokes equations that describes viscous, incompressible creeping flow. In particular, we consider the lid-driven cavity case where a square cavity consisting of three rigid walls with no-slip conditions and a lid moving with a tangential unit velocity. We consider that the uncertainty comes from the viscosity term and aim to predict the velocity field. See Section B.1 for details.
4.3 Evaluation metrics
We use the following three metrics to evaluate the performance of different methods. For the approximations of the solution $u(m)$, we compute the relative error in the $L^{2}(\Omega)$ norm and $H^{1}(\Omega)$ norm on the test dataset $\mathcal{D}_{\text{test}}=\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\text{test}}}$, that is,
where
For the approximation of the Jacobian $du(m;\cdot)$, we compute the relative error (for the discrete Jacobian) in the Frobenius norm on $\mathcal{D}_{\text{test}}$ along random directions $\omega=\{\omega_{i}\}_{i=1}^{N_{\text{dir}}}$, that is,
$\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{m^{(i)}\in\mathcal{D}_{\text{test}}}\frac{\|d\hat{u}(m^{(i)};\omega)-du(m^{(i)};\omega)\|_{F}}{\|du(m^{(i)};\omega)\|_{F}},$
where $\omega_{i}$ are samples drawn from the same distribution as $m$.
4.4 Main results
We compare the prediction errors measured in the above three evaluation metrics and computational cost in data generation and neural network training for four neural operator architectures, including DeepONet [1], FNO [2], DINO [20], and our DE-DeepONet. We also add experiments to demonstrate the performance of the FNO trained with the derivative loss (DE-FNO) and the DeepONet trained with input dimension reduction but without the derivative loss (DeepONet (ASM) or DeepONet (KLE)). For FNO, we use additional position embedding [4] that improves its approximation accuracy in our test cases. For DINO, we use ASM basis (v.s. KLE basis) as the input reduced basis and POD basis (v.s. output ASM basis [20]) as the output reduced basis, which gives the best approximation accuracy. For the input reduced basis of DE-DeepONet, we also test and present the results for both KLE basis and ASM basis.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Mean relative errors ($\pm$ standard deviation) over 5 random seeds of neural network training for a varying number of training samples for the [top: hyperelasticity; bottom: Navier–Stokes] equation using different methods. Relative errors in the $L^{2}(\Omega)$ norm (left) and $H^{1}({\Omega})$ norm (middle) for the prediction of $u=(u_{1},u_{2})$. Right: Relative error in the Frobenius (Fro) norm for the prediction of $du(m;\omega)=(du_{1}(m;\omega),du_{2}(m;\omega))$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Mean relative errors ($\pm$ standard deviation) over 5 random seeds versus model training time for the Navier–Stokes equations when the number of training samples is [top: 16; bottom: 256].
Test errors.
In Figure 1, we show a comparison of different methods for the hyperelasticity and Navier–Stokes problem in predicting the solution (in $L^{2}$ norm, left, and $H^{1}$ norm, middle) and its derivative (in Frobenius norm, right) with respect to parameter in the direction of $N_{\text{dir}}=128$ test samples $\{\omega_{i}\}_{i=1}^{N_{\text{dir}}}$. First, we can observe significant improvement of the approximation accuracy of DE-DeepONet compared to the vanilla DeepONet for all three metrics in all cases of training samples. Second, we can see that FNO leads to larger approximation errors than DINO and DE-DeepONet for all metrics except when number of training samples is 1024. We believe that the reason why FNO performs better than DE-DeepONet and DINO when training samples are large enough is mainly due to the use of input dimensionality reduction in DE-DeepONet and DINO (where the linear reduction error cannot be eliminated by increasing training samples) whereas in FNO we use full inputs. We also see that DE-FNO performs the best among all models when the training samples are sufficient (256 or 1024), although in the compensation of much longer training time shown in Figure 2. Third, we can see that the approximation accuracy of DINO is similar to DE-DeepONet (ASM) but requires much longer inference time as shown in Table 5 (see Section 5 and Section B.3.2 for reasons). In DINO, the output reduced basis dimension is set to be smaller than or equal to the number of training samples as the output POD basis are computed from these samples, i.e., 16 for 16 samples and 64 for $\geq 64$ samples. Increasing the output dimension beyond 64 does not lead to smaller errors in our test. Finally, we can observe that DE-DeepONet using ASM basis leads to smaller errors than using KLE basis, especially for the Navier-Stokes problem.
Table 1: Output reconstruction error with 16 input reduced bases
Dataset Relative $L^{2}$ error
KLE ASM
Hyperelasticity 3.8% 2.7%
Navier–Stokes 17.4% 5.8%
Moreover, we present the output reconstruction error due to the input dimension reduction using KLE or ASM basis in  Table 1. The errors provide the lower bound of the relative errors in $L^{2}(\Omega)$ norm of DINO and DE-DeepONet. We can see that using the ASM basis results in a lower output reconstruction error than the KLE basis (more significant difference observed in the more nonlinear Navier–Stokes equations). See Section B.6 for the decay of the reconstruction error with increasing number of reduced basis.
In addition, we provide full visualizations of the ground truth and prediction of both solution and derivatives in Section B.7. Visually, DE-DeepONet (ASM) consistently provides the best estimate (in terms of pattern and smoothness similarity with the ground truth) for both the solution and its derivative when the training data is limited (16 or 64 samples).
Data generation computational cost.
We use MPI and the finite element library FEniCS [31] to distribute the computational load of offline data generation to 64 processes for the PDE models considered in this work. See Table 2 for the wall clock time in generating the samples of Gaussian random fields (GRFs), solving the PDEs, computing the $r=16$ reduced basis functions (KLE or ASM) corresponding to the $16$ largest eigenvalues, generating the derivative labels, respectively. In Table 4, We also provide the total wall clock time of data generation of DE-DeepONet (ASM) (we only includes the major parts – computing high fidelity solution, ASM basis and dm labels [16 directions]) when $N_{\text{train}}=16,64,256,1024$ using 16 CPU processors.
Table 2: Wall clock time (in seconds) for data generation on 2 $\times$ AMD EPYC 7543 32-Core Processors
Dataset Process
GRFs PDEs KLE ASM dm labels
$(N_{\text{all}}=2000)$ $(N_{\text{all}}=2000)$ $(r=16)$ $(r=16)$ $(N_{\text{all}}=2000,r=16)$
($64$ procs) ($64$ procs) ($1$ procs) ($16$ procs) ($64$ procs)
Hyperelasticity $1.1$ $9.7$ $0.4$ $1.4$ $19.5$
Navier–Stokes $1.9$ $99.1$ $1.3$ $9.7$ $125.5$
Table 3: Wall clock time (in seconds) for data generation with different number of training samples using 16 CPU processors
$\text{Dataset}\textbackslash N_{\text{train}}$ PDEs + ASM basis + dm labels
16 64 256 1024
Hyperelasticity 2 5 16 61
Navier–Stokes 17 38 124 470
Table 4: Wall clock time (seconds/iteration with batch size 8) for training on a single NVIDIA RTX A6000 GPU
Dataset Model
DeepONet FNO DINO DE-FNO DE-DeepONet
Hyperelasticity 0.007 0.015 0.007 0.215 0.022
Navier–Stokes 0.007 0.015 0.007 0.216 0.033
Table 5: Total wall clock time (in seconds) for each model inferring on 500 test samples of both the solution and dm in 128 random directions, using a single GPU and a single CPU (except where specified)
Dataset Model
DeepONet FNO/DE-FNO DINO1 DE-DeepONet Numerical solver
1 GPU + 1 CPU/16 CPUs 0 GPU + 16 CPUs
Hyperelasticity 3 33 69/7 10 166
Navier–Stokes 3 33 2152/151 18 1103
1 The inference time of DINO is dominated by the time required to compute evaluations of all finite element basis functions at the grid points using FEniCS (which may not be the most efficient, see Section B.3.2). Even though these grid points overlap with parts of the finite element nodes—allowing us to skip evaluations by extracting the relevant nodes—for a fairer comparison with DE-DeepONet (in terms of its ability to evaluate at any arbitrary point), we assume they are arbitrary points requiring explicit evaluation.
Model training computational cost.
We present comparisons of the wall clock time of each optimization iteration (with batch size 8) of different methods in Table 4 and convergence plot (error versus training time) in Figure 2 and the figures in Section B.5. We find that incorporating derivative loss leads to longer training time as expected. However, when the training data are limited, the increased computation cost is compensated for a significant reduction of errors. We note that there are potential ways to further reduce the training cost, e.g., by training the model with additional derivative loss only during the later stage of training, or by using fewer points for computing the derivative losses in each iteration. Additionally, thanks to the dimension reduction of the input, we can define a relatively small neural network and thus are able to efficiently compute the derivatives using automatic differentiation.
5 Related work
Our work is related to Sobolev training for neural networks [32], which was found to be effective in their application to model distillation/compression and meta-optimization. In the domain of surrogate models for parametric partial differential equations, our work is more closely related to derivative-informed neural operator (DINO) [20] which is based on a derivative-informed projected neural network (DIPNet) [17], and presents an extension to enhance the performance of the DeepONet. Compared to DINO, although the DeepONet architecture (and its formulation of dm loss) requires longer training time, it offers the following advantages: (1) Potentially shorter inference time. The additional trunk net (which receives spatial coordinates) allows us to quickly query the sensitivity of output function at any point when input function is perturbed in any direction. While DINO can only provide the derivative of the output coefficients respect to the input coefficients (we call reduced dm), in order to compute the sensitivity at a batch of points, we need to post process the reduced dm by querying the finite element basis on these points and computing large matrix multiplications; (2) Greater flexibility and potential for improvements. Although both DeepONet and DINO approximate solution by a linear combination of a small set of functions, these functions together in DeepONet is essentially the trunk net, which is "optimized" via model training, whereas in DINO, they are POD or derivative-informed basis precomputed on training samples. When using DINO, if we encounter a case where the training samples not enough to accurately compute the output basis, the large approximation error between the linear subspace and solution manifold will greatly restrict the model prediction accuracy (see Figure 2 when $N_{\text{train}}=16$). And the reduced dm labels only supports linear reduction of output. However, it is possible that we can further improve DeepONet by, e.g., adding physical losses (to enhance generalization performance) and Fourier feature embeddings (to learn high-frequency components more effectively) on the trunk net [5] and replacing the inner product of the outputs of two networks by more flexible operations [33, 9] (to enhance expressive power). The dm loss formulation of our work is broadly suitable any network architecture that has multiple subnetworks, where at least one of them receives high-dimensional inputs.
6 Discussion
In this work, we proposed a new neural operator–Derivative-enhanced Deep Operator Network (DE-DeepONet) to address the limited accuracy of DeepONet in both function and derivative approximations. Specifically, DE-DeepONet employs a derivative-informed reduced representation of input function and incorporates additional loss into the loss function for the supervised learning of the derivative of the output with respect to the inputs of the branch net. Our experiments for nonlinear PDE problems with high variations in both input and output functions demonstrate that adding this loss term to the loss function greatly enhances the accuracy of both function and derivative approximations, especially when the training data are limited. We also demonstrate that the use of derivative loss can be extended to enhance other neural operators, such as the Fourier neural operator.
We presented matrix-free computation of the derivative label and the derivative-informed dimension reduction for a general form of PDE problems by using randomized algorithms and linearized PDE solves. Thanks to this scalable approach, the computational cost in generating the derivative label data is shown to be only marginally higher than generating the input-output function pairs for the test problems, especially for the more complex Navier–Stokes equations which require more iterations in the nonlinear solves than the hyperelasticity equation.
Limitations:
We require the derivative information in the training and dimension reduction using ASM, which may not be available if the explicit form of the PDE is unknown or if the simulation only provides input-output pairs from some legacy code. Another limitation is that dimension reduction of the input function plays a key role in scalable data generation and training, which may not be feasible or accurate for intrinsically very high-dimensional problems such as high frequency wave equations. Such problems are also very challenging and remain unsolved by other methods to our knowledge.
Acknowledgements
We would like to thank Jinwoo Go, Dingcheng Luo and Lianghao Cao for insightful discussions and helpful feedback.
Appendix A Proofs
A.1 Proof of  Theorem 1
@@
We first show how to compute $p$ given the PDE residual $\mathcal{R}(m,u)=0$. Since $u$ is uniquely determined by $m$, we can write $u=u(m)$ so that $\mathcal{R}(m,u(m))\equiv 0$ holds for any $m\in V^{\text{in}}$. Thus, for any $\varepsilon>0$ and $\psi\in V^{\text{in}}$, we have
@@
where $v$ is a test function in $V^{\text{out}}$.
@@
Let $q:=(\partial_{u}\mathcal{R})^{-*}p$. We can solve for $q$ via the weak form
$\langle(\partial_{u}\mathcal{R})^{*}q,v\rangle=\langle p,v\rangle,$
or equivalently,
$\langle q,(\partial_{u}\mathcal{R})v\rangle=\langle p,v\rangle,$ (19)
where $v$ is a test function in $V^{\text{out}}$.
By Equation 18, we have $w=-(\partial_{m}\mathcal{R})^{*}q$. For any test function $v\in V^{\text{in}}$, it holds that
@@
By Equation 9, we use $N_{\text{grad}}$ samples $\{(m^{(i)},u^{(i)})\}_{i=1}^{N_{\text{grad}}}$ to compute the Monte Carlo estimate of the action of operator $\mathcal{H}$ on any function $\psi\in V^{\text{in}}$, that is,
Remark.
@@
When using the double pass randomized algorithm to obtain the first $r$ eigenpairs in  Equation 10, we need to compute the action of $\mathcal{H}$ on $2(r+s)$ random functions in $V^{\text{in}}$ (e.g., their nodal values are sampled from the standard Gaussian distribution), where $s\in\mathbb{N}^{+}$ is typically a small oversampling parameter, often chosen between $5$ and $20$. To speed up the computation, we first compute the LU factorization of the matrices resulting from the discretization of the linear PDEs in Equation 17 and Equation 19. Then the action of $d^{*}u(m^{(i)};du(m^{(i)};\cdot))$ on these random functions can be efficiently computed via the forward and backward substitution. Furthermore, the computational time can be significantly reduced by parallelizing the computation of the average value in Equation 21 across multiple processors.
A.2 Proof of the equivalence of Equation 6 and Equation 7